### v1

# PreprintToPaper dataset: connecting bioRxiv preprints with journal publications

Abstract
The PreprintToPaper dataset connects bioRxiv preprints with their corresponding journal publications, enabling large-scale analysis of the preprint-to-publication process. It comprises metadata for 145,517 preprints from two periods, 2016–2018 (pre-pandemic) and 2020–2022 (pandemic), retrieved via the bioRxiv and Crossref APIs. Each record includes bibliographic information such as titles, abstracts, authors, institutions, submission dates, licenses, and subject categories, alongside enriched publication metadata including journal names, publication dates, author lists, and further information. Preprints are categorized into three groups: Published (formally linked to a journal article), Preprint Only (unpublished), and Gray Zone (potentially published but unlinked). To enhance reliability, title and author similarity scores were calculated, and a human-annotated subset of 299 records was created for evaluation of Gray Zone cases. The dataset supports diverse applications, including studies of scholarly communication, open science policies, bibliometric tool development, and natural language processing research on textual changes between preprints and their published versions. The dataset is publicly available in CSV format via Zenodo.
Background & Summary
Preprints have become an essential channel for the rapid dissemination of scientific findings, particularly in the life sciences, where timely access to results can accelerate discovery and collaboration. Their importance was intensified during the COVID-19 pandemic, when preprints served as a primary means of sharing the newest research across domains before formal peer review [1, 2].
The PreprintToPaper dataset [3], which we describe in this paper, links bioRxiv preprints with their subsequent journal publications, allowing for large-scale analysis of the preprint-to-publication process. It includes metadata on more than 145,000 preprints from two distinct periods (2016–2018, the pre-pandemic period, and 2020–2022, the COVID-19 pandemic period), with information on titles, authors, abstracts, institutions, licenses, submission and publication dates, and subject categories. Preprints are categorized into “Published”, “Preprint Only”, and “Gray Zone” cases, based on automated linking procedures combining DOI information, title similarity, and author similarity. In addition to the main dataset, a human-annotated subset of gray zone cases is provided to support evaluation of the automated matching approach  [^1] This subset includes records with title similarity (title_match_score = 0.75) and was independently checked by two annotators. Further details on the annotation procedure is provided in the Methods section (Step 5: Gray Zone Verification)..
The motivation for creating the PreprintToPaper dataset is to provide a comprehensive resource for studying how preprints evolve into published journal articles—or remain unpublished [4, 5]. While some preprints are eventually linked to formal publications, many others stay as preprints without ever appearing in journals. This dataset is the first systematic effort to automatically collect and link metadata from bioRxiv [6] with publication records, making it possible to analyze patterns in the preprint-to-publication process at scale [7].
We identify the following potential reuse cases for the PreprintToPaper dataset.
  • Policy and evaluation studies: Provides evidence for assessing the role of preprints in accelerating scientific dissemination, informing open science policies, and tracking uptake during events such as the COVID-19 pandemic [2, 8, 9]. To enable comparison, we included preprints from 2016-2018, which we designate as the pre-pandemic period.
  • Scholarly communication research: Enables analysis of publication delays, changes in titles, abstracts, and authorship between preprints and journal articles, and differences across research fields or time periods [10, 11].
  • Tool and method development: Serves as a reference dataset for testing algorithms in informetrics, metadata linking, and similarity matching (e.g., title and author disambiguation).
  • Natural language processing and text analysis: Supports research on linguistic and structural changes between preprints and their published versions, such as title reformulations, abstract rewriting, and shifts in author order or contribution statements.
In the following, we summarize related work that has used preprint data from bioRxiv.
Research on bioRxiv has shown that articles posted as preprints tend to receive more citations and online attention than comparable articles not deposited, although the causal mechanisms behind this effect remain uncertain [12]. Surveys of authors indicate that the main motivations for posting preprints are to increase visibility and accelerate dissemination, while concerns about peer review and lack of awareness are the main barriers [2]. Large-scale bibliometric analyses further confirm a citation and altmetric advantage for bioRxiv-linked publications, with preprints themselves increasingly cited and shared on social media [12]. During the COVID-19 pandemic, preprints on bioRxiv and medRxiv played an unprecedented role, being accessed, cited, and shared at higher rates, and influencing both journalistic reporting and policy discussions [1]. An analysis of biomedical publishing during the first months of the COVID-19 pandemic reveals a sharp rise in publication volume, faster acceptance times for COVID-19 papers, and reduced international collaboration, with these shifts occurring largely at the expense of non-COVID-19 research [8]. Other studies of bioRxiv preprints find that around 30% remain unpublished, most are posted too late to enable feedback, and nearly half of published preprints are concentrated in four major publishers, suggesting the platform is used more for priority setting and visibility than for prepublication peer review [4, 13].
Below, we provide an overview of related datasets based on bioRxiv preprints and collections of COVID-19–related preprints.
The Rxivist database provides a snapshot of preprints from bioRxiv and medRxiv collected via a custom web crawler, enabling readers to sort, filter, and analyze tens of thousands of preprints based on a PostgreSQL-importable database [14, 7].
The covid19_preprints dataset compiled weekly updated metadata on COVID-19–related preprints from multiple sources (Crossref, DataCite, arXiv, and RePEc), using keyword-based matching and deduplication procedures to track their distribution over time [15, 1].
The PreprintToPaper dataset offers two main advantages over the covid19_preprints dataset and the Rxivist database:
  • It stores both initial and final versions of preprints, as well as detailed metadata (title, abstract, authors, journal, publication date) about their published versions.
  • It identifies articles that are not formally linked but may have been published (Gray Zone cases).
Find a detailed comparison of the three datasets in Table 1.
Table 1: Comparison of the covid19_preprints, Rxivist, and PreprintToPaper datasets
Aspect covid19_preprints Rxivist PreprintToPaper (Ours)
Time span 2017–2021 Nov 2013 – Feb 2023 2016–2018 (pre-pandemic) and 2020–2022 (pandemic)
Coverage Multiple preprint servers (bioRxiv, arXiv, OSF, ESSOAR, Figshare, Zenodo, etc.) bioRxiv and medRxiv bioRxiv
Scope COVID-19 related preprints only All subjects (bioRxiv + medRxiv) All subjects (bioRxiv)
Size 66,130 records 223,541 records 145,517 records
Preprint information Source, identifier (DOI), posted date, title, abstract Title, abstract, DOI, category, submission date, authors, institutions Title, abstract (initial + last), DOI, category, submission date (initial + last), authors (initial + last), corresponding authors, institutions, license, version number, category, fulltext JATS XML
Publication information None Published DOI, journal name Published DOI, title, abstract, authors, journal name, publication date (from Crossref)
Versioning Only one version per preprint Only one version per preprint Both the first and last versions at the relevant period
Publication linking None Published DOIs partially incomplete (crawling stopped early) Official bioRxiv links + enriched with Crossref + additional Gray Zone investigation ($\approx$19,000 potential unlinked publications)
Methods
Dataset generation workflow
Figure 1 illustrates the data generation workflow for the PreprintToPaper dataset. The process consists of six main steps: (1) collecting metadata from the bioRxiv preprint platform; (2) retrieving additional metadata from Crossref based on the published DOIs listed in the bioRxiv records; (3) categorizing preprints into published and preprint-only; (4) identifying unlinked publications in bioRxiv (Gray Zone); (5) verifying Gray Zone preprints using author-based metrics (author match score and author count difference); and (6) generating the final dataset.
Figure 1: PreprintToPaper dataset creation workflow. In Step 5 (Gray Zone verification), a human-annotated subset of 299 rows was created, limited to cases with a title match score of 0.75 and covering both periods (2016–2018 and 2020–2022). Two annotators manually checked the corresponding abstracts.
Step 1: bioRxiv Metadata Retrieval
The first step in this study was to obtain metadata from the bioRxiv preprint platform using the official bioRxiv API [16]. The API has the following structure: server indicates the source of the data (e.g., bioRxiv or MedRxiv), interval indicates the desired date range, cursor indicates the pagination of the results, and format indicates the output format (e.g., JSON). The following data fields were collected through the API, and the retrieved metadata fields were mapped to the following structure in the dataset (see Table 2).
Table 2: Mapping of bioRxiv API fields to the dataset columns
bioRxiv API field Dataset column Description
DOI biorxiv_doi Unique digital identifier of the preprint
Title biorxiv_title Title of the preprint
Authors biorxiv_authors List of all authors
Corresponding author biorxiv_author_corresponding Name of the corresponding author
Corresponding author institution biorxiv_author_corresponding_institution Institution of the corresponding author
Date biorxiv_submission_date Submission date of the preprint
Version biorxiv_version Version number of the preprint
Type biorxiv_type Document type of the preprint
License biorxiv_license License
Category biorxiv_category Subject category
JATS XML path biorxiv_jatsxml Link to the XML file
Abstract biorxiv_abstract Abstract text
Published biorxiv_published_doi Publication DOI in a journal
Data were collected for two periods: 2016–2018 (pre-pandemic period) and 2020–2022 (COVID-19 pandemic period). The collected metadata was then processed. Because most preprints have multiple versions, only records that met the following criteria were retained during the processing stage:
  1. 1.
    If the preprint is in the relevant period and both the first and last versions of the preprint are available, then both versions were kept.
  2. 2.
    If the preprint has only one version and this version is the first, then this version was also retained.
If only non-initial versions were available during the relevant period, the corresponding records were deleted. Additionally, a pivoting step was applied to the entire dataset to prevent DOI duplication. To avoid this, each preprint was stored only once, but non-repeating fields were stored in separate columns. For example, the title, authors, corresponding author, submission dates, and abstract of both the first and last versions were stored in new columns. For this purpose, the _1st and _last suffixes were systematically added to the column names (e.g., biorxiv_title_1st and biorxiv_title_last).
This filtering aims to track changes between the first and last versions of the preprints, as well as changes that occurred when they were published in journals.
Step 2: Crossref Metadata Retrieval
For published preprints, the journal DOI is stored in the bioRxiv dataset. About two weeks after journal publication, bioRxiv automatically adds a link to the published article [17]. Detailed metadata for these articles was retrieved via the official Crossref API [18], by replacing the doi field with the Digital Object Identifier of the corresponding article. This query returns an extensive metadata package for the published version. The following fields were retrieved from the Crossref database (see Table 3).
Table 3: Mapping of Crossref API fields to the dataset columns
Crossref API field Dataset column Description
Container-title crossref_journal_name Name of the journal in which the article was published
Title crossref_title Official title of the published version
Author crossref_authors List of authors
Published-online crossref_online_publication_date Date the article was first published online
Published-print crossref_issue_online_date Date associated with the volume or issue of the journal in which the article appeared
The official publication dates of 500 published articles were manually compared with the online_publication_date and issue_online_date obtained from Crossref. In most cases, the online_publication_date matched with the actual publication date of the article.
Based on the results of manual checks, we retained the online_publication_date when available and used the issue_online_date otherwise. Both values were then merged into a single column, crossref_publication_date. Additionally, a new column, crossref_publication_type, was created to indicate whether the date refers to the online or issue version.
Step 3: Categorization of Preprints
In the third step, the dataset was categorized based on whether the preprints had DOIs linked to published articles on the bioRxiv preprint server.
  1. 1.
    If a published DOI was available for the preprint, the preprint was included in the published group.
  2. 2.
    If a DOI was not available, the preprint was categorized as preprint only.
The categorization result was stored in the column custom_status.
Step 4: Identification of unlinked publications in bioRxiv - Gray Zone
Although bioRxiv metadata is expected to include DOIs for all published preprints, verification revealed that in some cases the DOI of the published version is missing, even when the preprint has been published in a journal. For example, the fourth version of a preprint with DOI 10.1101/170381 was submitted to bioRxiv on August 6, 2020, but the same article was later published in Brain Structure and Function with DOI 10.1007/s00429-020-02136-0 on September 18, 2020. Nevertheless, the preprint still appears as preprint only in the bioRxiv record.
Previous studies have also documented this issue [7, 1, 5]. For example, Abdill and Blekhman (2019) found that 37.5% of 120 bioRxiv preprints were missing publication links. Fraser et al. (2021) reported that 7.6% of 12,788 preprints had missing publication links. Cabanac et al. (2021) reported that 60.3% of preprints marked as published on medRxiv and three other servers lacked corresponding publication links.
To identify such cases, we compared the titles of works marked as preprint only with the titles of published articles in the Crossref database. First, a separate request was sent to the Crossref database for each DOI, and matching titles were identified. Python’s SequenceMatcher library was then used to assess title similarity. The algorithm determines the longest common subsequence between two texts, accounts for the total length of matching parts, and calculates a similarity coefficient ranging from 0 to 1. A coefficient close to 1 indicates a complete match, whereas a value near 0 indicates substantial differences. Matches with a similarity index of $0.75$ or higher were classified into the Gray Zone category. This threshold was chosen because, in tests comparing preprints already marked as published in bioRxiv with their corresponding journal versions, results in this range proved most reliable in Figure 2.
Figure 2: Distribution of published papers by title match score
The calculated title similarity values were stored in the title_match_score column of the dataset. Additionally, for the preprints classified as Gray Zone, the corresponding metadata retrieved from the Crossref API including journal name, title, published date, author list, and publication date was stored in these fields crossref_journal_name, crossref_title, crossref_authors, crossref_publication_date, crossref_publication_type.
Step 5: Gray Zone Verification (Author Metrics)
To further improve the accuracy of preprints in the Gray Zone category, additional author-based metrics were applied. At this stage, the author names from the bioRxiv and Crossref metadata were compared. Two main metrics were calculated:
  • Author Count Difference – the difference between author counts in the bioRxiv and Crossref records.
  • Author Match Score – the similarity score between the author lists.
For the second metric, Unicode normalization was first performed to minimize differences in name format, diacritical marks were separated and removed, and all letters were reduced to lowercase. Then, for each record, the author list was first divided into individual authors by the “;” separator, and each author line was tokenized by splitting it into words using the space character. As a result, a nested (i.e., list-within-a-list) structure consisting of authors was constructed for each line. Based on this structure, the SequenceMatcher library was applied, and the author’s match score was calculated. This approach allowed for more accurate results in cases such as variations in the order of first and last names or when only initials were recorded.
Additionally, to validate these metrics, a human-annotated subset with a title match score of 0.75 was created, representing the lower bound of our title similarity threshold. Two annotators manually checked the corresponding bioRxiv and Crossref abstracts, and inter-annotator agreement was high.
Thus, the publication status of preprints in the Gray Zone category was determined more accurately using the author match score and author count difference metrics.
Step 6: Final Dataset Generation
In the final stage, several additional calculations were performed to enrich the dataset with further features. For example, the difference between the submission dates of the first and last versions of each preprint was calculated, as well as the difference between the official journal publication date and the last submission date on bioRxiv for published preprints. These values were stored in the custom_biorxivVersion_dateDifference and custom_submission&publication_dateDiff columns.
These operations made it possible to track the time dynamics between preprint versions and their subsequent journal publication. The finalized dataset was then saved in CSV format.
Data Record
The dataset is based on data collected from the bioRxiv [16] and Crossref [18] databases. It is stored in CSV format and is publicly available via Zenodo. The dataset consists of two main files:
  1. 1.
    The Main dataset includes metadata of preprints collected in 2016–2018 and 2020–2022. This file contains primary fields obtained from the bioRxiv API (DOI, title, authors, corresponding institutions, abstract, version, category, license, etc.) and additional fields obtained from the Crossref API (e.g., journal name, title in the journal, publication date, author list). In addition, there are custom fields calculated by us: title match score, author match score, author count difference, the difference between the first and last version submission dates, and the difference between the last preprint submission date and the publication date. Only fields that are not self-explanatory are described in detail below:
    • custom_status: Classification indicating the publication status of preprints: Published (officially published or confirmed with a DOI in bioRxiv), Preprint Only (unpublished), or Gray Zone (no official link in bioRxiv, but a possible publication identified based on title and author match).
    • biorxiv_published_doi: DOI of published versions of the preprint. This information is collected from two sources: (1) official publication DOIs provided by bioRxiv; (2) DOIs additionally identified in Gray Zone cases.
    • custom_submission&publication_dateDiff: The difference in days between the last preprint submission date and the publication date of the published version.
    • custom_biorxivVersion_dateDifference: The difference in days between the submission dates of the first and last bioRxiv versions.
    • biorxiv_version_last: The serial number of the last version of the preprint available on bioRxiv (e.g., a value of “5” indicates that five versions have been published).
    • biorxiv_jatsxml: A URL that references a JATS XML file containing the full text and metadata of the preprint, if available.
    • crossref_publication_date: The publication date retrieved from Crossref. This date can refer either to the online publication or the issue publication.
    • crossref_publication_date_type: Indicates the type of date given in the crossref_publication_date column: online publication or issue publication. If the online publication date is available, it is preferred; otherwise, the issue publication date is used.
    • author_count_diff: The difference between the number of authors in the preprint and the number of authors in the published version. Positive values indicate that the published version has more authors, while negative values indicate fewer.
    • title_match_score: The similarity score between the preprint and the published article titles. Values range from 0 to 1 and are calculated with the SequenceMatcher algorithm. Only matches within the 0.75–1.0 range were considered potential publication candidates.
    • author_match_score: Similarity score between the preprint and published article author lists. Ranges from 0 to 1 and is primarily used to verify Gray Zone cases, with higher values indicating stronger similarity.
    Table 4 presents the distribution of preprints in the dataset by period and category (Gray Zone, Preprint Only, Published) along with the total numbers.
    Period Gray Zone Preprint Only Published Total Preprints
    2016 359 1 019 3 343 4 721
    2017 841 2 314 8 191 11 346
    2018 1 121 4 115 12 943 18 179
    2020 3 746 8 884 26 081 38 711
    2021 5 292 8 987 22 539 36 818
    2022 7 731 10 494 17 517 35 742
    Total 19 090 35 813 90 614 145 517
    Table 4: Distribution of preprints in the PreprintToPaper dataset
  2. 2.
    The Human-Annotated Subset, containing 299 records, includes only cases with a title match score of $0.75$. Preprints from both periods (2016–2018 and 2020–2022) were selected, and two annotators verified potential matches by manually checking the corresponding abstracts from bioRxiv and Crossref. To support evaluation of the Gray Zone category, this file is provided as an additional resource. The dataset includes the following variables: year, biorxiv_doi, suspected_published_doi, author_match_score, annotator1, and annotator2.
    • suspected_published_doi: Suspected DOI of the published article. Includes DOIs identified for Gray Zone preprints where a potential published version was found but could not be fully confirmed.
    • annotator.1 / annotator2: Labels provided independently by two annotators to validate Gray Zone matches. Values: True (confirmed as correct match), False (not a valid match), or NA (uncertain).
Technical Validation
Figure 3: (a) Combined annotations versus author match score for records characterized by the title match score equal to 0.75. Symbols are data points, and the curve comes from the fit to the logistic regression model (1). Data points are artificially jittered around TRUE and FALSE values to avoid overlap. (b) The fraction of preprints that were eventually published as journal articles versus the year of preprint submission. Triangles represent data without the gray zone records, circles take into account gray zone records, with the author match score larger than 0.47 used to distinguish between preprints and eventually published journal papers.
To assess the validity of the proposed approach, all the records characterized by the $\textit{title\_match\_score}=0.75$ (in total 299 records) have been selected and annotated by two referees; each of them had an option to mark a record as TRUE (the identified journal publication was based on the preprint), FALSE (there was no connection between the preprint and the publication) or NA (the annotator was unable to make a call). The Cohen’s kappa coefficient [19] for non-NA data among annotators is equal to $\kappa=0.86$, indicating strong agreement among the annotators [20].
The annotations have been then merged by taking the conjunction of two annotations (i.e., if only if both annotators deemed the identified publication to be true, it was considered as such); if either of the annotators regarded the publication as NA it was assigned NA. The resulting distribution gives 69% of positive recognition and 28% of falsely assigned publications, with the remaining 3% that cannot be resolved. These outcomes suggest that the title matching score alone might be an insufficient indicator. Figure 3a presents the combined annotation versus author match score, presenting evidence that preprint – publication pairs with a low level of author match score are more likely to be among false positives than ones characterized by an author match score close to 1. To provide a more qualitative description of this fact and motivated by the observed relation between the combined annotation and author match score seen in Figure 3a, we used the logistic regression model[21] given by
$\log\left(\frac{\pi}{1-\pi}\right)=\beta_{0}+\beta_{1}x$ (1)
where $\pi$ is interpreted as the probability of the dependent variable (annotation) equaling a success, $x$ reflects author match score and $\beta_{0},\beta_{1}$ are regression coefficients. The model was fitted on the set of non-NA data, resulting in $\beta_{0}=-2.99\pm 0.46$ and $\beta_{1}=6.36\pm 0.67$ (see the curve in Figure 3a). The location parameter $\mu=-\beta_{0}/\beta_{1}$ – the midpoint of the curve, where $pi$ crosses 1/2 is $\mu=0.47\pm 0.12$ where the standard deviation of $\mu$ has been calculated using propagation of uncertainty[22], i.e., $\sigma^{2}_{\mu}=|\beta_{0}/\beta_{1}|\sqrt{\sigma^{2}_{\beta_{0}}/\beta^{2}_{0}+\sigma^{2}_{\beta_{1}}/\beta^{2}_{1}}$. To assess the validity of the logistic regression approach, we performed 10-fold cross-validation[21], obtaining the accuracy=0.92 and Cohen’s $\kappa=0.81$.
Let us now see the ramifications of the author match score threshold $\mu$ on the possible use of the dataset. Figure 3b presents the fraction of preprints that were eventually published as journal articles $f$ versus the year of preprint submission. The first series, marked by triangles, disregards the gray zone papers, taking into account only records marked as “published”, which leads to a spectacular decrease in $f$ by over 30% between years 2016 and 2022. However, if we take into account that all gray zone records with author match score exceeding $\mu=0.47$ can be regarded as published ones, the decrease does not exceed 11%.
Usage Notes
PreprintToPaper dataset can be used in a variety of studies: one of the principal ideas is connected to the issue of information overload[23] in academia – a notion that is widely understood[24] (e.g., paper tsunami[25]) but for which no measures have been defined yet. The very construction of our dataset (i.e., a set of preprints submitted before / during the outbreak of the COVID-19 pandemic), along with the information about the period between the preprint submission, can help find the patterns of preprint-to-paper transformations under the stress exerted by the pandemic[25]. In the same manner, the question of the characteristics of “unpublished preprints” – treated in a similar way as uncited papers[26, 27] – could be examined using the provided dataset. These include stylometric factors: title and abstract length, text complexity (e.g., Herdan’s measure[28]), readability scores (e.g., Gunning FOG index[29]) as well as text sentiment, number of authors, and other that have been shown to play role in the previous studies linking textual factors and paper popularity[30, 31, 32].
Code Availability
All custom scripts used in this study are publicly available in the GitHub repository: GitHub
The principal code was developed in Python 3.9+. The main external libraries used are pandas, requests, and numpy. Other libraries (unicodedata, re, difflib, datetime, time) are included in the Python standard library. All packages required for use are listed in the requirements.txt file.
The code consists of scripts for different stages: downloading metadata from the bioRxiv API (download_biorxiv.py), selecting the first and last version and creating the pivot structure (filter_1st_and_last_version.py, pivot_biorxiv_data.py), getting published metadata from the Crossref API (download_crossref.py), categorizing preprints (categorize_preprints.py), detecting “Gray Zone” cases (find_missing_links_crossref.py), verifying by author match (verify_missing_links_by_author_match.py), standardizing dates and calculating differences (standardize_crossref_dates.py, merge_publication_dates.py, calculate_date_difference.py).
The code for carrying out technical validation and for creating plots shown in Fig. 3 is developed in R 4.5.1[33] using the following external libraries (R packages): tidyverse[34], cowplot[35], psych[36], and caret[37]. The code consists of a single script technical_validation.R.
There are no restrictions on obtaining and using the code.
Acknowledgements
The authors of this work were funded by the European Union under the Horizon Europe grant OMINO – Overcoming Multilevel INformation Overload (grant number 101086321, http://ominoproject.eu). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. F.B. also acknowledges support from Deutsche Forschungsgemeinschaft (DFG) under grant number MA 3964/15-3, the SocioHub project. J.S. also acknowledges support from the Polish Ministry of Education and Science under the programme entitled International Co-Financed Projects and from a special grant for leaders of European projects coordinated by the Warsaw University of Technology.
Author contributions statement
P.M. conceived the idea, F.B. wrote the code, gathered the data, F.B. and J.S. annotated, validated, and analyzed the data. All authors drafted the first version of the manuscript.
Competing interests
The authors declare no competing interests.